# Probing AGN Feedback in Dwarf Galaxies with Spatially Resolved NIR Coronal Lines from JWST

### v1

@@ Abstract
We present the first spatially resolved investigation of near-infrared coronal lines in dwarf galaxies hosting active galactic nuclei (AGN), using JWST/NIRSpec integral field spectroscopy. Coronal lines (CLs), which are forbidden transitions from highly ionized species with ionization potentials up to 450 eV, act as sensitive tracers of the AGN ionizing continuum and feedback processes. Across four dwarf galaxies with ionized gas outflows traced by the optical [O III] lines, we report the detection of 16 unique species of near-infrared CLs. Line ratio diagnostics indicate that photoionization from the AGN dominates the excitation of CLs. We find that the coronal line region in dwarf galaxies, traced by the various CLs, extends up to 0.5 kpc, and can constitute up to 10% of their host galaxy size. Correlations between CL luminosities and [O III] ionized gas outflow properties are consistent with a scenario in which AGN-driven outflows likely facilitate the detection of CLs and contribute to their extent. Several CLs, including [Si VI], [Si VII], and [Mg VIII], exhibit a secondary broad component (W80 300 km s-1). If we interpret this spatially compact gas as part of an outflow, this would indicate that the outflowing gas includes a wide range of ionizations. The estimated energetics imply this highly ionized component is compact yet powerful enough to perturb gas in the central regions of the host dwarfs. These results indicate that AGN in low-mass galaxies may produce outflows capable of influencing their structure and evolution.
@@ 1 Introduction
Over the past decade, a growing number of studies have revealed that active galactic nuclei (AGN) are not limited to massive galaxies, but also occur in dwarfs with stellar masses below (2003ApJ...588L..13F; 2013ApJ...775..116R; 2014AJ....148..136M; 2018ApJ...863....1C; 2020ApJ...898L..30M; 2022ApJ...937....7S; 2025ApJ...982...10P). Their presence in such low-mass systems provides new constraints on black hole seed formation and the role of feedback in shallow gravitational potentials (2020ARA&A..58...27I; 2021NatRP...3..732V; 2021MNRAS.503.3568K; 2022MNRAS.516.2112K; 2023ApJ...946...51C). Traditionally, stellar feedback from supernovae and young stars has been regarded as the dominant regulator of star formation and gas loss in dwarfs, but recent observations show that AGN can also drive galaxy-wide outflows (2018MNRAS.476..979P; 2019ApJ...884...54M; 2024ApJ...965..152L; 2025ApJ...979...26S; 2025A&A...697A.235R). In some cases, the mass and kinetic energy outflow rates of AGN in dwarf galaxies are comparable to those in massive galaxies when scaled by AGN luminosity (2019ApJ...884...54M; 2020ApJ...905..166L), and AGN-driven winds have been shown to be faster and more energetic than stellar-driven outflows in hosts of similar mass and at similar redshifts (2023ApJ...950...33A).
| Name | Short name | Redshift | log(M∗/M⊙) | log(MBH/M⊙) | log[L(AGN)] | SFR | W |
| (erg s-1) | (M⊙ yr-1) | (km s-1) | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) |
| J084234.51+031930.7 | J0842 | 0.0291 | 9.3 | 5.8 0.5 (i) | 43.1 0.4 | 0.3 | 500 |
| J090613.75+561015.5 | J0906 | 0.0467 | 9.4 | 5.4 0.3 (ii) | 43.7 0.8 | 0.3 | 980 |
| J095418.16+471725.1 | J0954 | 0.0329 | 9.1 | 5.0 0.4 (ii) | 43.9 1.1 | 0.3 | 730 |
| J100935.66+265648.9 | J1009 | 0.0145 | 8.8 | 5.1 0.1 (iii) | 43.0 1.0 | 0.1 | 210 |
@@
Note. — Columns (1): SDSS name of the target. (2): Short name used in this paper. (3): Redshift, as measured from a combination of stellar absorption features and Pa 1.86 m. (4): Stellar mass from the NASA Sloan Atlas (2011AJ....142...31B) with an average error of 0.5 dex. (5): Black hole mass derived using broad line fits and Eddington luminosities from the following references: (i) 2019ApJ...884...54M, (ii) 2013ApJ...775..116R; 2017AA...602A..28M, (iii) 2014AJ....148..136M. We caution that the estimated masses are uncertain, by 0.5 dex or more. (6): Bolometric AGN luminosity based on extinction-corrected [O III] luminosity (2020ApJ...905..166L; 2024ApJ...965..152L). (7): Upper limit on the star formation rate (SFR) based on the extinction-corrected [O II] 3726, 3729 from 2020ApJ...905..166L, assuming that one-third of the [O II] emission is from the star formation activity, following 2005ApJ...629..680H. (8): Median W80 measurements of the broad, outflowing component in [O III] determined from KCWI data (2020ApJ...905..166L).
@@
Unlike stellar feedback, which mostly acts locally and is limited by the star formation rate, AGN can inject energy on galaxy-wide scales, often with outflow velocities that exceed the escape velocity of dwarfs (2019ApJ...884...54M). Even short-lived AGN episodes may therefore have disproportionate effects compared to continuous stellar activity, regulating star formation, removing gas and metals, and enriching the circumgalactic medium. If AGN feedback proves effective at these mass scales, it would imply that black hole growth and galaxy evolution are coupled across the full mass spectrum, challenging the paradigm that stellar feedback dominates in dwarfs while AGN feedback shapes massive galaxies (2011IAUS..277..273S). This raises a key question: can AGN feedback play an equally significant role in dwarf galaxy evolution as in their massive counterparts?
@@
A major obstacle in answering this question lies in confirming that the observed outflows are indeed AGN-driven. Common tracers of ionized gas such as [O III] 5007 can suffer contamination from stellar processes, making it challenging to distinguish between stellar- and AGN-powered winds in these low-mass systems. This necessitates the need for observational diagnostics that are both sensitive to AGN photoionization and largely immune to stellar contamination (2019ApJ...870L...2C).
@@
Coronal lines (CLs), which are forbidden transitions from highly ionized species with ionization potentials up to 450 eV, satisfy these requirements. Their excitation typically requires the hard radiation field from an accreting black hole. Although recent observations have detected CLs in stellar winds from massive stars (2024ApJ...976L..25L), the presence of highly ionized (150 eV) and extended CLs are unlikely to be produced by stellar populations alone (2021ApJ...906...35S; 2022ApJ...927..165R). Thus, CLs provide a unique window into the otherwise unobservable ionizing continuum of Type-2 AGN, and have been used to constrain the spectral energy distribution of the central engine, estimate black hole masses (2018ApJ...861..142C; 2022MNRAS.510.1010P), and trace the most highly ionized phases of AGN-driven outflows (2011ApJ...739...69M; 2018MNRAS.481L.105M; 2022ApJ...940L...5U; 2023MNRAS.524..143F; 2023ApJ...942L..37A; 2024ApJ...965..103B). Their high ionization potentials ensure that the emission is free from contamination by stellar outflows. In massive galaxies, CLs have been observed to extend beyond the narrow-line region, often tracing interactions between AGN jets and the interstellar medium (2006ApJ...653.1098R; 2021ApJ...920...62N; 2025ApJ...984..170M). Because CLs likely trace only the AGN-powered, most highly ionized phases of outflows, they provide a direct probe of AGN feedback. This is particularly critical in dwarf galaxies, where the shallow potential wells make them more susceptible to the gas-clearing effects of even modest AGN-driven winds (2017ApJ...839L..13S).
@@
Despite their potential to identify AGN and AGN-feedback in more massive galaxies, CLs in dwarf galaxies remain largely unexplored. Photoionization modeling predicts that intermediate-mass black holes (IMBHs) in dwarfs should produce stronger CL emission due to their harder ionizing spectral energy distribution (SED) (2018ApJ...861..142C). Yet, only a small number of dwarf galaxies have reported detections of CLs, both in the optical and near-infrared (2021ApJ...911...70B; 2021ApJ...912L...2C; 2021ApJ...922..155M; 2023ApJ...946L..38R; 2024ApJ...975...60A). Ground-based NIR spectroscopy of dwarfs hosting AGN-driven, ionized-gas outflows has revealed high-ionization CLs (2021ApJ...911...70B), but their spatial extents, kinematics, and connection to AGN feedback remain poorly constrained.
@@
In this paper, we use JWST/NIRSpec integral field spectroscopy to spatially map the coronal line region (CLR) in dwarf galaxies hosting AGN and to examine its relationship to AGN-driven outflows. Our resampled 0.05 pixel scale of NIRSpec allows us to resolve the inner structure of the CLR, probe the spatial extent of NIR CLs, including several transitions at wavelengths 2.4 m, which are difficult to access from the ground, and confirm the AGN origin of observed outflows. By directly linking CL emission to feedback processes, we provide new constraints on the physical conditions, energetics, and influence of AGN in the low-mass regime. The paper is organized as follows: in Section 2.2, we present our JWST observations and data reduction; Section 3 describes our analysis and spatial mapping techniques; Section 4 presents our CL detections and the ionizing source; Section 5 discusses the coronal line region and implications for AGN feedback in dwarf galaxies; and Section 6 summarizes our findings. Throughout, we assume a flat CDM cosmology with = 70 km s-1 Mpc-1, = 0.3, and = 0.7.
2 Observations and Data reduction
2.1 Sample selection: Presence of ionized gas outflows
@@
We began with 29 Baldwin–Phillips–Terlevich (BPT; 1981PASP...93....5B)-selected, spatially extended, AGN-hosting dwarf galaxies (2013ApJ...775..116R; 2014AJ....148..136M; 2019ApJ...884...54M). Follow-up Keck/LRIS long-slit spectroscopy and Keck/KCWI IFU data isolated eight systems with galaxy-scale warm-ionized outflows traced by [O III] doublet (2019ApJ...884...54M; 2020ApJ...905..166L) (Figure 1). Near-IR spectroscopy with Keck/NIRES then detected CLs in five of these eight (2021ApJ...911...70B). From this set, we selected four targets listed in Table 1 with stellar masses 109.5 M⊙ for JWST/NIRSpec IFU that sample a range in stellar mass. Further details of the JWST sample will be provided in T. Bohn et al. 2025, in prep.
@@
The [O III] emission profile in these galaxies has a broad, often blueshifted wing. 2019ApJ...884...54M; 2020ApJ...905..166L model this with an additional Gaussian distinct from the narrow systemic gas with median bulk velocities of and W80 up to . HST/COS observations of three of the fastest systems (2024ApJ...965..152L) show definite outflows, producing blueshifted UV absorption in the C II, C IV, Si II, and Si IV lines. Photoionization modeling for one of the galaxies, J0954, places the absorbing gas at 0.5 kpc, suggesting a galactic-scale impact. The measured kinetic-energy outflow rates also exceed those of starburst winds at comparable SFRs (2024ApJ...965..152L). Keck/NIRES detections of [Si VI] (and other NIR CLs) often show higher outflow velocities than [O III], and the lack of J-band CN absorption features argues against a dominant young-stellar origin (2021ApJ...911...70B). The lack of [O III] 4363 detection makes it challenging to estimate direct gas-phase metallicities for the observed targets. However, since BPT-selected AGNs in dwarf galaxies are expected to be biased toward higher metallicities (2013ApJ...775..116R), we assume that the metallicities are nearly solar.
2.2 JWST/NIRSpec Observations and Data Reduction
@@
The data were obtained from the JWST Cycle 2 program (PID 3663, P.I.: Thomas Bohn). All four targets were observed with NIRSpec IFS (2022A&A...661A..80J; 2023PASP..135c8001B). J0906 was observed on 17 March 2024, J1009 on 9 May 2024, J0954 on 8 March 2024 and J0842 on 7 April 2024. Data were collected using the NIRSpec IFU high resolution mode in the G140H/F100LP, G235H/F170LP, and G395H/F290LP grating/filter combinations. The resulting wavelength coverage obtained with our choice of grating/filter combination was 0.9 - 5.2 m with a nominal resolving power of R 2700 (110 km s-1). We obtained dedicated background observations for the four targets instead of the standard LeakCal exposures in order to subtract the background to detect any faint CL.
@@
We downloaded the raw data files from the Barbara Mikulski Archive for Space Telescopes (MAST) and subsequently processed them with the JWST Science Calibration pipeline version 1.15.1 (2023zndo...7577320B) under the Calibration Reference Data System (CRDS) context jwst_1253.pmap. The level 1 files downloaded from MAST were processed using the Detector1 stage (stage 1), which performs detector-level corrections and generates count-rate images. In order to remove artifacts generated by cosmic-ray hits on the detector, we applied the snowball flagging for the jump step during the first stage of the pipeline. We used the dedicated background exposures to perform the pixel-to-pixel background subtraction for each target in stage 2. The stage2 images were then resampled and combined into a final data cube through the Calwebb_spec3 processing (stage3) using the ‘drizzle’ method with a spaxel size of 0.05 to enhance the spatial resolution. The resulting data cubes display the well-known sinusoidal modulations in the spectrum caused by the limitation of the pipeline to correctly trace the target across the IFU slices during the rectification process. To correct these artifacts, also known as “wiggles”, we applied a custom code similar to that developed by 2023A&A...679A..89P that masks out the detected lines and models and subtracts the sinusoidal variations (see W. Matzko et al. 2025, in prep). After completing the pipeline reduction, we removed any remaining pixel outliers from the data cubes using the methods outlined in 2024PASP..136d4503H.
3 Analysis
3.1 Emission line fitting and spatial maps generation

### v2

@@ Abstract
We present the first spatially resolved investigation of near-infrared coronal lines in dwarf galaxies hosting active galactic nuclei (AGN), using JWST/NIRSpec integral field spectroscopy. Coronal lines (CLs), which are forbidden transitions from highly ionized species with ionization potentials up to 450 eV, act as sensitive tracers of the AGN ionizing continuum and feedback processes. Across four dwarf galaxies with ionized gas outflows traced by the optical [O III] lines, we report the detection of 16 unique species of near-infrared CLs. Line ratio diagnostics indicate that photoionization from the AGN dominates the excitation of CLs. We find that the coronal line region in dwarf galaxies, traced by the various CLs, extends up to 0.5 kpc, and can constitute up to 10% of their host galaxy size. Correlations between CL luminosities and [O III] ionized gas outflow properties are consistent with a scenario in which AGN-driven outflows likely facilitate the detection of CLs and contribute to their extent. Several CLs, including [Si VI], [Si VII], and [Mg VIII], exhibit a secondary broad component with W80 (the line width enclosing 80% of the total flux) 300 km s-1. If we interpret this spatially compact gas as part of an outflow, this would indicate that the outflowing gas includes a wide range of ionizations. The estimated energetics imply this highly ionized component is compact yet powerful enough to perturb gas in the central regions of the host dwarfs. These results indicate that AGN in low-mass galaxies may produce outflows capable of influencing their structure and evolution.
@@ I Introduction
Over the past decade, a growing number of studies have revealed that active galactic nuclei (AGN) are not limited to massive galaxies, but also occur in dwarfs with stellar masses below (Filippenko and Ho, 2003; Reines et al., 2013; Moran et al., 2014; Chilingarian et al., 2018; Mezcua and Domínguez Sánchez, 2020; Salehirad et al., 2022; Pucha et al., 2025). Their presence in such low-mass systems provides new constraints on black hole seed formation and the role of feedback in shallow gravitational potentials (Inayoshi et al., 2020; Volonteri et al., 2021; Koudmani et al., 2021, 2022; Chadayammuri et al., 2023). Traditionally, stellar feedback from supernovae and young stars has been regarded as the dominant regulator of star formation and gas loss in dwarfs, but recent observations show that AGN can also drive galaxy-wide outflows (Penny et al., 2018; Manzano-King et al., 2019; Liu et al., 2024; Salehirad et al., 2025; Rodríguez Morales et al., 2025). In some cases, the mass and kinetic energy outflow rates of AGN in dwarf galaxies are comparable to those in massive galaxies when scaled by AGN luminosity (Manzano-King et al., 2019; Liu et al., 2020), and AGN-driven winds have been shown to be faster and more energetic than stellar-driven outflows in hosts of similar mass and at similar redshifts (Aravindan et al., 2023).
| Name | Short name | Redshift | log(M∗/M⊙) | log(MBH/M⊙) | log[L(AGN)] | SFR | v | W |
| (erg s-1) | (M⊙ yr-1) | (km s-1) | (km s-1) | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) |
| J084234.51+031930.7 | J0842 | 0.0291 | 9.3 | 5.8 0.5 (i) | 43.1 0.4 | 0.3 | -110 | 500 |
| J090613.75+561015.5 | J0906 | 0.0467 | 9.4 | 5.4 0.3 (ii) | 43.7 0.8 | 0.3 | -70 | 980 |
| J095418.16+471725.1 | J0954 | 0.0329 | 9.1 | 5.0 0.4 (ii) | 43.9 1.1 | 0.3 | -60 | 730 |
| J100935.66+265648.9 | J1009 | 0.0145 | 8.8 | 5.1 0.1 (iii) | 43.0 1.0 | 0.1 | -20 | 210 |
@@
Note. — Columns (1): SDSS name of the target. (2): Short name used in this paper. (3): Redshift, as measured from a combination of stellar absorption features and Pa 1.86 m. (4): Stellar mass from the NASA Sloan Atlas (Blanton et al., 2011) with an average error of 0.5 dex. (5): Black hole mass derived using broad line fits and Eddington luminosities from the following references: (i) Manzano-King et al. 2019, (ii) Reines et al. 2013; Marleau et al. 2017, (iii) Moran et al. 2014. We caution that the estimated masses are uncertain, by 0.5 dex or more. (6): Bolometric AGN luminosity based on extinction-corrected [O III] luminosity (Liu et al., 2020, 2024). (7): Upper limit on the star formation rate (SFR) based on the extinction-corrected [O II] 3726, 3729 from Liu et al. 2020, assuming that one-third of the [O II] emission is from the star formation activity, following Ho 2005. (8): Median v50 (velocity at the 50th percentile of the total velocity) measurements of the broad, outflowing component in [O III] determined from KCWI data (Liu et al., 2020). (9): Median W80 (line width that includes 80% of the total flux)) measurements of the broad, outflowing component in [O III] also determined from KCWI data (Liu et al., 2020).
@@
Unlike stellar feedback, which mostly acts locally and is limited by the star formation rate, AGN can inject energy on galaxy-wide scales, often with outflow velocities that exceed the escape velocity of dwarfs (Manzano-King et al., 2019). Even short-lived AGN episodes may therefore have disproportionate effects compared to continuous stellar activity, regulating star formation, removing gas and metals, and enriching the circumgalactic medium. If AGN feedback proves effective at these mass scales, it would imply that black hole growth and galaxy evolution are coupled across the full mass spectrum, challenging the paradigm that stellar feedback dominates in dwarfs while AGN feedback shapes massive galaxies (Silk, 2011). This raises a key question: can AGN feedback play an equally significant role in dwarf galaxy evolution as in their massive counterparts?
@@
A major obstacle in answering this question lies in confirming that the observed outflows are indeed AGN-driven. Common tracers of ionized gas such as [O III] 5007 can suffer contamination from stellar processes, making it challenging to distinguish between stellar- and AGN-powered winds in these low-mass systems. This necessitates the need for observational diagnostics that are both sensitive to AGN photoionization and largely immune to stellar contamination (Cann et al., 2019).
@@
Coronal lines (CLs), which are forbidden transitions from highly ionized species with ionization potentials up to 450 eV, satisfy these requirements. Their excitation typically requires the hard radiation field from an accreting black hole. Although recent observations have detected CLs in stellar winds from massive stars (Law et al., 2024), the presence of highly ionized (150 eV) and extended CLs are unlikely to be produced by stellar populations alone (Satyapal et al., 2021; Richardson et al., 2022). Thus, CLs provide a unique window into the otherwise unobservable ionizing continuum of Type-2 AGN, and have been used to constrain the spectral energy distribution of the central engine, estimate black hole masses (Cann et al., 2018; Prieto et al., 2022), and trace the most highly ionized phases of AGN-driven outflows (Müller-Sánchez et al., 2011; May et al., 2018; U et al., 2022; Fonseca-Faria et al., 2023; Armus et al., 2023; Bianchin et al., 2024). Their high ionization potentials ensure that the emission is free from contamination by stellar outflows. In massive galaxies, CLs have been observed to extend beyond the narrow-line region, often tracing interactions between AGN jets and the interstellar medium (Rodríguez-Ardila et al., 2006; Negus et al., 2021; Matzko et al., 2025). Because CLs likely trace only the AGN-powered, most highly ionized phases of outflows, they provide a direct probe of AGN feedback. This is particularly critical in dwarf galaxies, where the shallow potential wells make them more susceptible to the gas-clearing effects of even modest AGN-driven winds (Silk, 2017).
@@
Despite their potential to identify AGN and AGN-feedback in more massive galaxies, CLs in dwarf galaxies remain largely unexplored. Photoionization modeling predicts that intermediate-mass black holes (IMBHs) in dwarfs should produce stronger CL emission due to their harder ionizing spectral energy distribution (SED) (Cann et al., 2018). Yet, only a small number of dwarf galaxies have reported detections of CLs, both in the optical and near-infrared (Bohn et al., 2021; Cann et al., 2021; Molina et al., 2021; Reefe et al., 2023; Aravindan et al., 2024). Ground-based NIR spectroscopy of dwarfs hosting AGN-driven, ionized-gas outflows has revealed high-ionization CLs (Bohn et al., 2021), but their spatial extents, kinematics, and connection to AGN feedback remain poorly constrained.
@@
In this paper, we use JWST/NIRSpec integral field spectroscopy to spatially map the coronal line region (CLR) in dwarf galaxies hosting AGN and to examine its relationship to AGN-driven outflows. Our resampled 0.05 pixel scale of NIRSpec allows us to resolve the inner structure of the CLR, probe the spatial extent of NIR CLs, including several transitions at wavelengths 2.4 m, which are difficult to access from the ground, and confirm the AGN origin of observed outflows. By directly linking CL emission to feedback processes, we provide new constraints on the physical conditions, energetics, and influence of AGN in the low-mass regime. The paper is organized as follows: in Section II.2, we present our JWST observations and data reduction; Section III describes our analysis and spatial mapping techniques; Section IV presents our CL detections and the ionizing source; Section V discusses the coronal line region and implications for AGN feedback in dwarf galaxies; and Section VI summarizes our findings. Throughout, we assume a flat CDM cosmology with = 70 km s-1 Mpc-1, = 0.3, and = 0.7.
II Observations and Data reduction
II.1 Sample selection: Presence of ionized gas outflows
@@
We began with 29 Baldwin–Phillips–Terlevich (BPT; Baldwin et al. 1981)-selected, spatially extended, AGN-hosting dwarf galaxies (Reines et al., 2013; Moran et al., 2014; Manzano-King et al., 2019). Follow-up Keck/LRIS long-slit spectroscopy and Keck/KCWI IFU data isolated eight systems with galaxy-scale warm-ionized outflows traced by [O III] doublet (Manzano-King et al., 2019; Liu et al., 2020) (Figure 1). Near-IR spectroscopy with Keck/NIRES then detected CLs in five of these eight (Bohn et al., 2021). From this set, we selected four targets listed in Table 1 with stellar masses 109.5 M⊙ for JWST/NIRSpec IFU that sample a range in stellar mass. Further details of the JWST sample will be provided in T. Bohn et al. 2025, in prep.
@@
The [O III] emission profile in these galaxies has a broad, often blue-shifted wing. Manzano-King et al. 2019; Liu et al. 2020 model this with an additional Gaussian distinct from the narrow systemic gas with median bulk velocities of and W80 up to . HST/COS observations of three of the fastest systems (Liu et al., 2024) show definite outflows, producing blue-shifted UV absorption in the C II, C IV, Si II, and Si IV lines. Photoionization modeling for one of the galaxies, J0954, places the absorbing gas at 0.5 kpc, suggesting a galactic-scale impact. The measured kinetic-energy outflow rates also exceed those of starburst winds at comparable SFRs (Liu et al., 2024). Keck/NIRES detections of [Si VI] (and other NIR CLs) often show higher outflow velocities than [O III], and the lack of J-band CN absorption features argues against a dominant young-stellar origin (Bohn et al., 2021). The presence of the AGN makes it challenging to measure direct (Te-based) gas-phase metallicities for our targets. A systematic analysis of AGN-hosting dwarf galaxies in the Reines et al. 2013 sample, which includes the targets studied in this paper, finds a mean metallicity of 0.3 Z⊙ for the sample as a whole (Cai et al., 2020), although individual galaxy values are not reported. Additionally, BPT-selected AGNs in dwarf galaxies are expected to be biased toward higher metallicities (Reines et al., 2013). The relatively high stellar masses of the galaxies in this sample make extremely low metallicities (0.01 Z⊙) unlikely. Throughout this work, we therefore adopt near-solar metallicities, consistent with prior studies of this sample (Liu et al., 2020; Bohn et al., 2021), and we also explore trends assuming sub-solar values (0.5 Z⊙ and 0.25 Z⊙).
II.2 JWST/NIRSpec Observations and Data Reduction
@@
The data were obtained from the JWST Cycle 2 program (PID 3663, P.I.: Thomas Bohn). All four targets were observed with NIRSpec IFS (Jakobsen et al., 2022; Böker et al., 2023). J0906 was observed on 17 March 2024, J1009 on 9 May 2024, J0954 on 8 March 2024 and J0842 on 7 April 2024. Data were collected using the NIRSpec IFU high resolution mode in the G140H/F100LP, G235H/F170LP, and G395H/F290LP grating/filter combinations. The resulting wavelength coverage obtained with our choice of grating/filter combination was 0.9 - 5.2 m with a nominal resolving power of R 2700 (110 km s-1). We obtained dedicated background observations for the four targets instead of the standard LeakCal exposures in order to subtract the background to detect any faint CL.
@@
We downloaded the raw data files from the Barbara Mikulski Archive for Space Telescopes (MAST) and subsequently processed them with the JWST Science Calibration pipeline version 1.15.1 (Bushouse et al., 2023) under the Calibration Reference Data System (CRDS) context jwst_1253.pmap. The level 1 files downloaded from MAST were processed using the Detector1 stage (stage 1), which performs detector-level corrections and generates count-rate images. In order to remove artifacts generated by cosmic-ray hits on the detector, we applied the snowball flagging for the jump step during the first stage of the pipeline. We used the dedicated background exposures to perform the pixel-to-pixel background subtraction for each target in stage 2. The stage2 images were then resampled and combined into a final data cube through the Calwebb_spec3 processing (stage3) using the ‘drizzle’ method with a spaxel size of 0.05 to enhance the spatial resolution. The resulting data cubes display the well-known sinusoidal modulations in the spectrum caused by the limitation of the pipeline to correctly trace the target across the IFU slices during the rectification process. To correct these artifacts, also known as “wiggles”, we applied a custom code similar to that developed by Perna et al. 2023 that masks out the detected lines and models and subtracts the sinusoidal variations (see W. Matzko et al. 2025, in prep). After completing the pipeline reduction, we removed any remaining pixel outliers from the data cubes using the methods outlined in Hutchison et al. 2024.
III Analysis
III.1 Emission line fitting and spatial maps generation